Simplify the expression
i^15

Answer by stanbon(75887)

Simplify the expression
i^15
---
Since i^4 is 1
i^8 and i^12 are also 1.
--
i^15 = i^12*i^3 = 1*(-i) = -i
===================================
Cheers,
Stan H.
=========